OpenStudy (anonymous):

how do you solve 6/(x-5)=4(x-3)?

OpenStudy (anonymous):

multiply both sides by (x-5) and post what you get.

OpenStudy (anonymous):

it's 6/(x-5)=4/(x-3)

OpenStudy (anonymous):

oh I see. so if a = b then 1/a = 1/b so take the reciprocals on both sides and post what you get.